# Emergence and Localisation of Semantic Role Circuits in LLMs

### v1

@@ Abstract
Despite displaying semantic competence, large language models’ internal mechanisms that ground abstract semantic structure remain insufficiently characterised. We propose a method integrating role-cross minimal pairs, temporal emergence analysis, and cross-model comparison to study how LLMs implement semantic roles. Our analysis uncovers: (i) highly concentrated circuits (89–94% attribution within $\leq$28 nodes); (ii) gradual structural refinement rather than phase transitions, with larger models sometimes bypassing localised circuits; and (iii) moderate cross-scale conservation (24–59% component overlap) alongside high spectral similarity. These findings suggest that LLMs form compact, causally isolated mechanisms for abstract semantic structure, and these mechanisms exhibit partial transfer across scales and architectures.
@@ 1 Introduction
Do LLMs develop abstract, causally functional representations of semantic structure? Large language models (LLMs) exhibit localised circuits for factual recall (goldowskydill2023localizing; meng2022locating), arithmetic (conmy2023towards; Stolfo2023AMI), and logical reasoning (kim2025reasoning). However, it remains unclear whether such mechanisms extend to the abstract relational semantic structure that underlies natural language understanding.
Current mechanistic studies often focus on specialised algorithmic behaviours in trained models (e.g., induction heads, copying, factual associations (meng2022locating; gpt2_attention_saes)), leaving two central gaps in our understanding of semantic representations in LLMs.
First, abstract semantic structure. Semantic roles (e.g., Agent, Theme, Instrument) formalise predicate–argument abstractions that generalise across surface forms and syntactic realisations (fillmore1976frame). Formally, a semantic role $r$ associates a predicate $p$ and argument position $i$ with a thematic relation: for instance, in “The children played in the garden”, the Location role links garden to the playing event whether expressed as “in the garden”, “at the garden”, or “outside in the garden” (see Appendix A for a fuller formalisation). From a mechanistic perspective, such predicate–argument binding requires integrating information across tokens and abstracting away from local surface cues, unlike behaviours explained by positional heuristics or lexical templates. Whether LLMs implement this binding through causally functional circuits thus remains an open question.
Second, temporal emergence. Most circuit analyses examine only final checkpoints, obscuring when semantic mechanisms arise, stabilise, and become computationally indispensable. Given evidence that several behaviours emerge through sharp transitions (e.g., in-context learning, algorithmic generalisation (wei2022emergent; nanda2023progress; he2024learning)), it is unclear whether semantic-role circuits likewise appear abruptly or gradually consolidate. Understanding this timeline is important both for training-time interventions (aljaafari2025trace; cheng2024potentiallimitationsllmscapturing) and for explaining how syntactic and semantic abstractions co-develop.
width=⟧figures/updated_Compass.png
@@
To address these questions, we introduce COMPASS (Compositional Predicate–Argument Semantic Structure; Fig. 1), a causal–temporal methodology that combines edge-attribution patching (hanna2024have) with training-time circuit tracking [^1] Supporting code and datasets are available at a public repository << anonymised url >>. It studies predicate–argument binding using role-cross minimal pairs, disentangles the formation of circuit structure from its functional engagement, and evaluates the transferability of role circuits across model scales. We apply COMPASS to the Pythia (14M–1B) and LLaMA-1B model families to answer:
  • ⟦leftmargin=*, itemsep=1pt, parsep=0pt, topsep=3pt⟧
@@
Across four roles (Beneficiary, Instrument, Location and Time), we find that role-binding circuits localise to compact sets of attention heads and MLPs, and that larger models reuse similar components while wiring them differently. The results also indicate that predicate–argument binding emerges through gradual refinement rather than sudden reorganisation. Circuit presence does not guarantee immediate functional use: models appear to allocate semantic capacity early and exploit it only later. By demonstrating causally determined, functional circuits for shallow semantic structure, our method provides mechanistic evidence that LLMs acquire structured semantic representations and offers new avenues for targeted circuit editing and training-time intervention.
@@
We summarise our contributions as: (i) introduce COMPASS, a causal–temporal method for discovering and tracking semantic-role circuits in LLMs; (ii) show that semantic-role information concentrates in small sets of components whose structural organisation stabilises before becoming functionally indispensable; and (iii) demonstrate partial cross-scale and cross-architecture transfer of these circuits between Pythia (14M–1B) and LLaMA-1B.
@@ 2 Related Work
Mechanistic Interpretability and Circuit Discovery.
Mechanistic interpretability seeks to reverse-engineer the computations of neural networks (bereska2024mechanistic). Foundational work analysed how attention heads, MLPs, and residual streams implement algorithms (elhage2021mathematical), leading to discoveries such as induction heads (olsson2022context) and circuits for task-specific behaviours (wang2022interpretability; conmy2023towards). Causal intervention methods form the core of modern circuit analysis: activation patching tests causal effects by swapping activations (meng2022locating), while gradient-based variants such as Attribution Patching (AtP) (nanda2023attribution; syed-etal-2024-attribution) improve scalability but suffer from gradient saturation. $\text{AtP}^{*}$ (kramar2024atp) introduces architectural fixes, including gradient dropout and QK decomposition. More recently, Edge Attribution Patching (EAP) (hanna2024have) attributes causal influence to individual attention edges, and EAP-IG mitigates saturation effects via Integrated Gradients, offering high faithfulness for fine-grained circuit analysis.
Parallel lines of work seek to disentangle polysemantic features. Sparse Autoencoders (SAEs) extract monosemantic latent features from superimposed representations (templeton2024scaling; bricken2023monosemanticity), with applications to scripts and relational knowledge (deng2025sparse). Transcoders (dunefsky2024transcoders) approximate component-level transformations via sparse bottlenecks, and attribution graphs (ameisen2025circuit) visualise feature interactions using surrogate models. However, surrogate-based approaches introduce reconstruction errors and do not guarantee causal necessity (gao2025scaling; kantamneni2025sparse). For studying the emergence of semantic circuits across training, methods that operate directly on the original model are required.
@@
Transformers often acquire new capabilities through sharp transitions in training dynamics, including grokking-style shifts from memorisation to generalisation (power2021grokking; nanda2023progress; aljaafari2025trace) and phase changes in in-context learning (wei2022emergent). Recent work also shows that linguistic structure develops progressively: subspaces associated with syntax and semantics become more coherent over training (muller-eberstein-etal-2023-subspace), and task-specific circuits such as IOI or subject–verb agreement emerge in coordinated phases (NEURIPS2024_47c7edad). However, most such analyses rely on probing or scalar metrics, leaving open whether these transitions reflect the emergence of causally functional circuits. To the best of our knowledge, no prior work combines causal circuit localisation with systematic temporal tracking of compositional semantic representations across training and model scale.
@@
While transformers achieve strong Semantic Role Labelling (SRL), classifying predicates and argument roles within sentences (chen2025semanticrolelabelingsystematical), most work examines their accuracy rather than internal mechanisms. Probing studies suggest a hierarchical organisation of linguistic knowledge, lexical features in lower layers, syntax in mid layers, semantics in higher layers (tenney2019bert; hewitt2019structural). However, probing reveals only linear separability, not causal involvement (caucheteux2021disentangling; conia-navigli-2022-probing). Circuit-level interpretability has mapped concrete behaviours such as IOI (wang2022interpretability), but has not addressed whether transformers implement causally functional circuits for abstract predicate–argument relations, motivating our focus on semantic-role mechanisms.
@@
Probing scales efficiently but lacks causal grounding; intervention methods such as path patching (GoldowskyDill2023LocalizingMB) and causal scrubbing (chan2022causal) provide strong guarantees but become computationally prohibitive across many checkpoints. Surrogate-based approaches (SAEs, transcoders) introduce reconstruction artefacts that hinder temporal and cross models comparisons. We therefore employ EAP-IG (hanna2024have), which provides path-specific causal attribution directly on the original model. We extends EAP-IG to a temporal setting, enabling causal analysis of when semantic-role circuits form, stabilise, and become functionally engaged across training and scale. A detailed comparison of interpretability methods is provided in Appendix G.
@@ 3.1 Semantic Role Circuits
To investigate predicate–argument binding, we adopt a role-conditioned continuation task grounded in frame semantics (fillmore1976frame) and PropBank-style role inventories (PropBank). The core idea is to vary the role-indicating scaffold (e.g., a prepositional phrase) while keeping the predicate and argument fillers fixed. If models represent abstract semantic roles, this behaviour should be mediated by localised role circuits $\mathcal{C}^{(r)}$ that map role-indicating scaffolds to role-appropriate continuations. We therefore construct role-cross minimal pairs that differ only in the scaffold, enabling isolation of role-specific behaviour. For example:
@@
Pairs use single-token role fillers and enforce token parity ($|\text{toks}(x^{(r)})|=|\text{toks}(x^{(s)})|$) to allow us to construct genuine minimal pairs. We additionally retain only items for which the model predicts the correct continuation in both variants, ensuring that any discovered circuits reflect functionally active behaviour (Appendix B).
@@
Task performance uses next-token accuracy:
$\begin{split}\text{CNP-Acc}=\\ \mathbb{E}_{(x^{(r)},y^{(r)})}\left\llbracket\mathds{1}\left\llbracket y^{(r)}=\operatorname*{arg\,max}_{v}P_{\theta}(v\mid x^{(r)})\right\rrbracket\right\rrbracket\end{split}$ (1)
Circuit attribution uses the negative log-probability of the role-appropriate target as the loss:
$\mathcal{L}_{\text{CNP}}(x^{(r)},y^{(r)})=-\log P_{\theta}(y^{(r)}\mid x^{(r)}),$ (2)
and EAP-IG attributes this loss to edges to identify components supporting role-appropriate predictions.
@@ 3.2 COMPASS: Causal-Temporal Circuit Discovery
COMPASS integrates causal localisation with temporal tracking to recover role circuits $\mathcal{C}^{(r)}$ that exhibit (i)causal necessity, ablating $\mathcal{C}^{(r)}$ reduces task performance; (ii) structural sparsity, $|\mathcal{C}^{(r)}|\ll|\mathcal{E}|$; and (iii) temporal stability, persistent structure after emergence. It also identifies emergence times $(\hat{t}_{c},t_{\text{cons}})$ characterising when circuits become functional. The procedure has three phases (full details in Appendix E).
@@
For each role pair $(x^{(r)},x^{(s)})$, EAP-IG (hanna2024have) computes edge scores by approximating their causal contribution to $\mathcal{L}_{\text{CNP}}$:
$S_{u\to v}^{\text{IG}}=\Delta_{u}^{\top}\bar{\mathbf{g}}_{v},\quad\text{where }\Delta_{u}=\mathbf{z}_{u}^{(r)}-\mathbf{z}_{u}^{(s)},$ (3)
and $\bar{\mathbf{g}}_{v}$ averages gradients along an IG interpolation path with $\alpha_{k}=k/m$ (m=5m{=}5). We normalise by total mass:
$\tilde{S}_{u\to v}^{\text{IG}}=S_{u\to v}^{\text{IG}}/\sum_{e\in\mathcal{E}}|S_{e}^{\text{IG}}|$ (4)
and extract the top-K=200K{=}200 edges by $|\tilde{S}_{u\to v}^{\text{IG}}|$ at each checkpoint $t$ to define $\mathcal{C}_{t}^{(r)}$. This threshold balances faithfulness saturation with interpretability. Node importance is induced from incident edge mass:
$\text{Importance}^{(r)}_{t}(\ell,h)=\sum_{\begin{subarray}{c}(u\to v)\in\mathcal{C}_{t}^{(r)}\\ v\in(\ell,h)\end{subarray}}|\tilde{S}_{u\to v}^{\text{IG}}|.$ (5)
Full EAP-IG implementation details appear in Appendix C.
@@
At each checkpoint $t$, we compute causal and structural signals: Causal signal: Faithfulness (mueller2025mib) measures how indispensable a circuit is:
$\text{Faithfulness}_{t}(\mathcal{C})=\frac{M_{t}(\mathcal{C})-M_{t}(\emptyset)}{M_{t}(\mathcal{E})-M_{t}(\emptyset)},$ (6)
where $M_{t}(\mathcal{C})$ is CNP accuracy when only edges in $\mathcal{C}$ are active (others zeroed), $M_{t}(\mathcal{E})$ is full-model performance, and $M_{t}(\emptyset)$ is the null baseline.
Structural signals. Circuit persistence is measured via Jaccard similarity across consecutive checkpoints:
$\text{Stability}_{t}(\mathcal{C})=\frac{|\mathcal{C}_{t}\cap\mathcal{C}_{t+\delta}|}{|\mathcal{C}_{t}\cup\mathcal{C}_{t+\delta}|},$ (7)
along with Top-$K$ node mass and Gini coefficient to quantify sparsification, detailed ablation protocols are in Appendix C.4.
@@
We define two temporal markers:
Consolidation time $t_{\text{cons}}$: first step where the top-$K$ node set stabilises:
$t_{\text{cons}}=\min\{t:\text{Stability}_{t}(\mathcal{C})\geq 0.6\text{ for }\geq 2\text{ steps}\}.$ (8)
Functional transition $\hat{t}_{c}$: change-point in faithfulness, estimated by the best-fitting two-segment linear model to $\{\text{Faithfulness}_{t}(\mathcal{C}_{t})\}_{t=0}^{T}$ via least-squares regression. The change-point maximising explained variance is:
$\hat{t}_{c}=\operatorname*{arg\,max}_{t}\,R^{2}\left(\text{PiecewiseLinear}(t)\right),$ (9)
with 95% confidence intervals via bootstrap resampling (n=1,000n=1{,}000). Change-point estimation details and sensitivity analyses appear in Appendix C.4.

### v2

@@ Abstract
Despite displaying semantic competence, large language models’ internal mechanisms that ground abstract semantic structure remain insufficiently characterised. To investigate whether and how LLMs develop causally functional representations of semantic roles, we introduce a causal-temporal methodology combining contrastive minimal pairs, edge-attribution circuit discovery, and training-time tracking. Our analysis reveals that LLMs encode semantic roles through highly localised circuits (89–92% attribution within $\leq$28 nodes) that emerge gradually via structural refinement rather than phase transitions. These circuits exhibit moderate cross-scale conservation (24–51% component overlap) alongside high spectral similarity, with larger models reusing similar components while rewiring connections. These findings suggest that LLMs form compact, causally isolated mechanisms for abstract semantic structure that exhibit partial transfer across scales and architectures.
@@ 1 Introduction
Do LLMs develop abstract, causally functional representations of semantic structure? Large language models (LLMs) exhibit localised circuits for factual recall (goldowskydill2023localizing; meng2022locating), arithmetic (conmy2023towards; Stolfo2023AMI), and logical reasoning (kim2025reasoning). However, it remains unclear whether such mechanisms extend to the abstract relational semantic structure that underlies natural language understanding. Current mechanistic studies often focus on specialised algorithmic behaviours in trained models (e.g., induction heads, copying, factual associations (meng2022locating; gpt2_attention_saes)), leaving two central gaps in our understanding of semantic representations in LLMs.
First, abstract semantic structure. Semantic roles (e.g., Agent, Theme, Instrument) constitute a prevalent descriptive component of predicate–argument structure that generalises across surface forms and syntactic realisations in natural language (fillmore1976frame). Formally, a semantic role $r$ associates a predicate $p$ and argument position $i$ with a thematic relation: for instance, in “The children played in the garden”, the Location role links garden to the playing event whether expressed as “in the garden”, “at the garden”, or “outside in the garden” (see App. A for a complete formalisation). Mechanistically, such predicate–argument binding requires integrating information across tokens and abstracting from surface cues, unlike behaviours explained by positional heuristics or lexical templates. Whether LLMs implement this binding through causally functional circuits however remains an open question.
Second, temporal emergence. Most circuit analyses examine only final checkpoints, obscuring when semantic mechanisms arise, stabilise, and become computationally indispensable. Given that several behaviours emerge through sharp transitions (e.g., in-context learning, algorithmic generalisation (wei2022emergent; nanda2023progress; he2024learning)), it is unclear whether semantic-role circuits likewise appear abruptly or gradually consolidate. Understanding this timeline is important both for training-time interventions (aljaafari2025trace2; cheng2024potentiallimitationsllmscapturing) and for explaining how syntactic and semantic abstractions co-develop.
@@
To address these questions, we introduce COMPASS (Compositional Predicate–Argument Semantic Structure; Fig.1), a causal–temporal methodology that combines edge-attribution patching (hanna2024have) with training-time circuit tracking [^1] Supporting code and datasets are available at a public repository << anonymised url >>. It investigates predicate–argument binding by isolating role-specific computation through contrastive prompts that differ only in their role-indicating scaffold (e.g., “delivered the package to the office” vs. “with the truck”), what we term role-cross minimal pairs, then tracks the causal pathways responsible for correct predictions. This approach disentangles the formation of circuit structure from its functional engagement and enables the evaluation of its transferability. We apply COMPASS across model scales and families to answer:
@@
We find that role-binding circuits localise to compact sets of attention heads and MLPs, with the results indicating that predicate–argument binding emerges through gradual refinement rather than sudden reorganisation. Circuit presence does not guarantee immediate functional use: models appear to allocate semantic capacity early and exploit it only later. Larger models reuse similar components while wiring them differently. By demonstrating causally determined, functional circuits for shallow semantic structure, our method provides mechanistic evidence that LLMs acquire structured semantic representations and offers new avenues for targeted circuit editing and training-time intervention.
@@
We summarise our contributions as: (i) introduce COMPASS, a causal–temporal method for discovering and tracking semantic-role circuits in LLMs; (ii) show that semantic-role information concentrates in small sets of components whose structural organisation stabilises before becoming functionally indispensable; and (iii) demonstrate partial cross-scale and cross-architecture transfer of these circuits.
@@ 2 Related Work
Mechanistic Interpretability (MI) and Circuit Discovery.
MI seeks to reverse-engineer neural network computations (bereska2024mechanistic). Foundational work analysed how attention heads, MLPs, and residual streams implement algorithms (elhage2021mathematical), leading to discoveries such as induction heads (olsson2022context) and task-specific circuits (wang2022interpretability; conmy2023towards). Causal intervention methods form the core of modern circuit analysis: activation patching tests causal effects by swapping activations (meng2022locating), while gradient-based variants such as Attribution Patching (AtP) (nanda2023attribution; syed-etal-2024-attribution) improve scalability but suffer from gradient saturation. $\text{AtP}^{*}$ (kramar2024atp) introduces architectural fixes; Edge Attribution Patching (EAP) (hanna2024have) attributes causal influence to individual edges, with EAP-IG mitigating saturation via Integrated Gradients. Parallel work on feature disentanglement uses Sparse Autoencoders (templeton2024scaling; bricken2023monosemanticity) and transcoders (dunefsky2024transcoders), but they are surrogate-based approaches that introduce reconstruction errors and do not guarantee causal necessity (gao2025scaling; kantamneni2025sparse).
@@
Transformers often acquire capabilities through sharp transitions, including grokking-style shifts (power2021grokking; nanda2023progress; aljaafari2025trace2) and phase changes in in-context learning (wei2022emergent). Recent work shows linguistic structure develops progressively: subspaces associated with syntax and semantics become more coherent over training (muller-eberstein-etal-2023-subspace), and task-specific circuits emerge in coordinated phases (NEURIPS2024_47c7edad). However, most analyses rely on probing or scalar metrics, leaving open whether transitions reflect emergence of causally functional circuits. To our knowledge, no prior work combines causal circuit localisation with systematic temporal tracking of compositional semantic representations across training and scale.
@@
While transformers achieve strong Semantic Role Labelling (chen2025semanticrolelabelingsystematical), most work examines accuracy rather than internal mechanisms. Probing studies suggest hierarchical linguistic knowledge organisation (tenney2019bert; hewitt2019structural), but probing reveals only linear separability, not causal involvement (caucheteux2021disentangling; conia-navigli-2022-probing). Circuit-level interpretability has mapped concrete behaviours such as IOI (wang2022interpretability) but has not addressed whether transformers implement causally functional circuits for abstract predicate–argument relations, motivating our focus on semantic-role mechanisms.
@@
Probing scales efficiently but lacks causal grounding; intervention methods such as path patching (GoldowskyDill2023LocalizingMB) and causal scrubbing (chan2022causal) provide strong guarantees but are computationally prohibitive across many checkpoints. Surrogate-based approaches introduce reconstruction artefacts hindering temporal comparisons. We employ EAP-IG (hanna2024have), which provides path-specific causal attribution directly on the original model, extending it to a temporal setting to analyse when semantic-role circuits form, stabilise, and become functionally engaged across training and scale (detailed comparison of methods in App. G).
@@ 3.1 Semantic Role Circuits
To investigate predicate–argument binding, we adopt a role-conditioned continuation task grounded in frame semantics (fillmore1976frame) and PropBank (PropBank). We vary the role-indicating scaffold while keeping predicate and argument fillers fixed. If models represent abstract semantic roles, this behaviour should be mediated by localised role circuits $\mathcal{C}^{(r)}$. We construct role-cross minimal pairs that differ only in the scaffold:
@@
Pairs use single-token fillers and enforce token parity ($|\text{toks}(x^{(r)})|=|\text{toks}(x^{(s)})|$). We retain only items where the model predicts correct continuations in both variants, ensuring circuits reflect functionally active behaviour (detailed in App. B).
@@
Task performance uses next-token accuracy: $\text{CNP}_{Acc}=\mathbb{E}_{(x^{(r)},y^{(r)})}\left[\mathds{1}\left[y^{(r)}=\operatorname*{arg\,max}_{v}P_{\theta}(v\mid x^{(r)})\right]\right].$ Circuit attribution uses the negative log-probability of the role-appropriate target: $\mathcal{L}_{\text{CNP}}(x^{(r)},y^{(r)})=-\log P_{\theta}(y^{(r)}\mid x^{(r)}),$ and EAP-IG attributes this loss to edges to identify components supporting role–appropriate predictions.
@@ 3.2 COMPASS: Causal-Temporal Circuit Discovery
COMPASS integrates causal localisation with temporal tracking to recover role circuits $\mathcal{C}^{(r)}$ that exhibit (i) causal necessity, (ii) structural sparsity ($|\mathcal{C}^{(r)}|\ll|\mathcal{E}|$), and (iii) temporal stability. It identifies emergence times $(\hat{t}_{c},t_{\text{cons}})$ characterising when circuits become functional. The procedure has three phases (full details of method and metrics are in App. E and D).
@@
For each role pair $(x^{(r)},x^{(s)})$, EAP-IG (hanna2024have) computes edge scores approximating their causal contribution to $\mathcal{L}_{\text{CNP}}$: $S_{u\to v}^{\text{IG}}=\Delta_{u}^{\top}\bar{\mathbf{g}}_{v},\quad$ where $\Delta_{u}=\mathbf{z}_{u}^{(r)}-\mathbf{z}_{u}^{(s)},$ and $\bar{\mathbf{g}}_{v}$ averages gradients along an IG interpolation path with $\alpha_{k}=k/m$ (we use m=5m{=}5). We normalise by total mass: $\tilde{S}_{u\to v}^{\text{IG}}=S_{u\to v}^{\text{IG}}/\sum_{e\in\mathcal{E}}|S_{e}^{\text{IG}}|$ and extract the top-K=200K{=}200 edges by $|\tilde{S}_{u\to v}^{\text{IG}}|$ at each checkpoint $t$ to define $\mathcal{C}_{t}^{(r)}$. Node importance is induced from incident edge mass: $\text{Importance}^{(r)}_{t}(\ell,h)=\sum_{\begin{subarray}{c}(u\to v)\in\mathcal{C}_{t}^{(r)}\\ v\in(\ell,h)\end{subarray}}|\tilde{S}_{u\to v}^{\text{IG}}|.$
@@
At each checkpoint $t$, we compute causal and structural signals, including metrics from mueller2025mib. Faithfulness measures circuit indispensability: $\text{Faithfulness}_{t}(\mathcal{C})={(M_{t}(\mathcal{C})-M_{t}(\emptyset))}/{(M_{t}(\mathcal{E})-M_{t}(\emptyset))},$ where $M_{t}(\mathcal{C})$ is $\text{CNP}_{acc}$ when only edges in $\mathcal{C}$ are active, $M_{t}(\mathcal{E})$ is full-model performance, and $M_{t}(\emptyset)$ is null baseline. Circuit persistence is measured via Jaccard similarity: $\text{Stability}_{t}(\mathcal{C})={|\mathcal{C}_{t}\cap\mathcal{C}_{t+\delta}|}/{|\mathcal{C}_{t}\cup\mathcal{C}_{t+\delta}|},$ along with Top-$K$ node mass and Gini coefficient.
@@
We define two temporal markers. Consolidation time $t_{\text{cons}}$ marks when the top-$K$ node set stabilises: $t_{\text{cons}}=\min\{t:\text{Stability}_{t}(\mathcal{C})\geq 0.6\text{ for }\geq 2\text{ steps}\}.$
Functional transition $\hat{t}_{c}$ is the change-point in faithfulness, estimated by the best-fitting two-segment linear model via least-squares regression: $\hat{t}_{c}=\operatorname*{arg\,max}_{t}\,R^{2}\left(\text{PiecewiseLinear}(t)\right),$
with 95% confidence intervals via bootstrap resampling (n=1,000n{=}1{,}000; details in App. D.4).